OpenStudy (anonymous):

Evaluate: |2.8 - k| + 3.4 if k = 8.1 A. 8.7 B. -1.9 C. 14.3

OpenStudy (anonymous):

A. 8.7

OpenStudy (anonymous):

\[\ \left| 2.8-3.4 \right|\] = 5.3 5.3+3.4= 8.7